OpenStudy (anonymous):

If a = 1.5, b = −2, and c = −5, then (a - b)^2 /c^2 is, rounded to the nearest hundredth, is equal to?

zepdrix (zepdrix):

Are you expected to do this problem without a calculator? :)

zepdrix (zepdrix):

Plugging in our values:\[\large\rm \frac{(a-b)^2}{c^2}\quad=\frac{(1.5--2)^2}{(-5)^2}\]

zepdrix (zepdrix):

Where you having trouble miss Diane? :)